# Benchmarking XAI Explanations with Human-Aligned Evaluations

### v1

@@ B.2 Dataset analysis
To further explore annotator preferences, we identified the top-12 XAI techniques selected by each annotator and visualized the results in the histogram shown in Figure 10. From this figure, we observe that classical methods such as LIME and SHAP stand out as the most frequently preferred. This suggests a strong preference for well-established saliency-based methods. Additionally, there is a notable inclination towards methods that probe the model’s reaction to input perturbations. A distinct aspect is the inclusion of B-cos, which generates explanations through the incorporation of a dedicated layer, offering a unique mechanism compared to other perturbation-based techniques. Of the 11 most popular techniques among annotators, only four are based on Concept Bottleneck Models (CBMs), indicating a general preference for saliency maps over concept-based explanations. Unlike Table 10, this analysis focuses on the top-12 techniques per annotator, removing the influence of votes among the top-12 techniques to reduce noise and better capture annotator preferences.
@@
Next, we aggregated the scores using majority voting and calculated QWC scores to measure the agreement between individual annotators and the aggregated score. We further analyzed the QWC scores by gender and age groups to assess any systematic differences in interpretation. As shown in Figure 12, the kappa scores indicate that there is generally consistent agreement across different age and sex groups, although older annotators show slightly less consistency. This highlights that while demographic factors may introduce some variation, they do not substantially impact the overall interpretability evaluation.
@@
We also investigated potential biases in the annotations themselves by examining the differences in how annotators approached CBM-based and saliency-based explanations. For CBM explanations, we focused on the text written by annotators in response to question Q0.1, assessing whether annotators preferred explanations that closely resembled their own textual responses. To quantify this, we transformed CBM explanations into text by concatenating the top concepts used in the explanation and calculated the BLEU (Papineni et al., 2002) and ROUGE scores (Lin, 2004) between these explanations and the annotators’ text responses. As shown in Figure 11, the ROUGE score reveals a slight correlation between the explanations and the annotators’ expectations for questions Q1, Q2, Q3, and Q4. This suggests that annotators are inclined to favor explanations that align with their preconceived notions, potentially introducing a bias toward consistency with their initial answers.
@@
Figure 11: Correlation between the various questions and key metrics for saliency and CBM explanations. For CBM, the criteria used are the BLEU (Papineni et al., 2002) and ROUGE scores (Lin, 2004) scores between the explanation and the text from question Q0.1. For saliency maps, the metrics include the total pixel sum (SUM_all), the sum of pixels within the annotator-provided bounding box (SUM_pos) from question Q0.2, the sum of pixels outside the bounding box (SUM_neg), and the entropy of the saliency map (Entropy).
Refer to caption Refer to caption
B.3 Additional results of the human evaluations
In Table 10, we present the average mode of votes for each XAI technique. the first observation is that it is difficult to observe clear differences among XAI methods. This is mainly due to the fact XAI methods are highly sensible to the backbone they are applied on, as noticed in Section 3.5. We observe also that the average score for saliency-based techniques across fidelity related questions is 2.47, while for CBMs, the average score is lower at 2.12. For saliency-based techniques across complexity related questions is 2.44, while for CBMs, the average score is also lower at 2.10. For saliency-based techniques across objectivity related questions is 2.46, while for CBMs, the average score is also lower at 2.11. For robustness related questions, the average scores are 3.59 for saliency-based techniques and 3.20 for CBMs.
Fidelity Complexity Objectivity Robustness
XAI Technique (Q1) (Q2-Q3) (Q4) (Q5-Q6)
GradCAM (ResNet50) 2.50 2.53 2.60 3.83
GradCAM (ViT-B) 2.77 2.86 3.00 3.57
GradCAM (CLIP-zero-shot) 1.86 1.78 1.76 3.90
LIME (ResNet50) 2.16 2.08 2.13 4.28
LIME (ViT-B) 3.06 2.98 2.99 3.89
LIME (CLIP-zero-shot) 2.48 2.49 2.51 4.25
SHAP (ResNet50) 2.63 2.66 2.65 3.74
SHAP (ViT-B) 2.87 2.85 2.89 3.64
SHAP (CLIP-zero-shot) 2.55 2.52 2.47 3.70
AblationCAM (ResNet50) 2.78 2.96 2.91 3.14
AblationCAM (ViT-B) 1.75 1.75 1.75 3.84
AblationCAM (CLIP-zero-shot) 1.27 1.34 1.36 3.64
EigenCAM (ResNet50) 2.23 2.23 2.41 3.21
EigenCAM (ViT-B) 3.31 3.25 3.15 3.39
EigenCAM (CLIP-zero-shot) 3.68 3.68 3.64 3.28
EigenGradCAM (ResNet50) 2.81 2.91 3.10 3.24
EigenGradCAM (ViT-B) 1.41 1.21 1.21 3.86
EigenGradCAM (CLIP-zero-shot) 2.29 2.10 2.05 3.67
FullGrad (ResNet50) 3.65 3.68 3.65 3.26
FullGrad (ViT-B) 1.65 1.50 1.65 4.00
GradCAMPlusPlus (ResNet50) 2.55 2.60 2.60 3.45
GradCAMPlusPlus (ViT-B) 2.19 2.00 1.90 3.93
GradCAMPlusPlus (CLIP-zero-shot) 1.80 1.88 1.85 4.13
GradCAMElementWise (ResNet50) 3.04 2.78 2.74 3.59
GradCAMElementWise (ViT-B) 1.47 1.53 1.53 3.90
GradCAMElementWise (CLIP-zero-shot) 2.39 2.33 2.35 3.24
HiResCAM (ResNet50) 2.96 2.77 2.83 3.52
HiResCAM (ViT-B) 1.45 1.30 1.30 4.10
HiResCAM (CLIP-zero-shot) 1.83 1.79 1.92 3.85
ScoreCAM (ResNet50) 2.68 2.55 2.45 3.36
ScoreCAM (ViT-B) 3.00 3.00 3.00 3.63
XGradCAM (ResNet50) 2.57 2.72 2.86 3.50
XGradCAM (ViT-B) 2.00 2.18 2.18 4.02
XGradCAM (CLIP-zero-shot) 2.50 2.31 2.21 4.34
DeepFeatureFactorization (ResNet50) 3.50 3.40 3.33 3.25
DeepFeatureFactorization (ViT-B) 2.46 2.66 2.69 3.45
DeepFeatureFactorization (CLIP-zero-shot) 2.94 2.92 3.22 3.61
BCos (ResNet50-BCos) 2.91 2.84 2.77 3.34
CLIP-QDA-sample 1.71 1.69 1.66 4.04
CLIP-Linear-sample 2.19 2.17 2.27 2.89
LIME_CBM (CLIP-QDA) 2.22 2.20 2.25 4.27
SHAP_CBM (CLIP-QDA) 2.44 2.33 2.29 3.81
LIME_CBM (CBM-classifier-logistic) 1.66 1.66 1.65 3.53
SHAP_CBM (CBM-classifier-logistic) 1.98 2.01 1.94 3.30
Xnesyl-Linear 1.72 1.77 1.77 3.82
RISE (CBM-classifier-logistic) 2.65 2.57 2.62 2.70
B.4 Evaluation Questions
@@
  • Q7: Could you rank the different qualities of the explanation in order of importance? Expectedness, Trustworthy, Complexity, Robustness, Objectivity.
@@
The results for Q7 are presented in Table 11. The data indicate that evaluators place a higher value on Trustworthiness and Complexity, while Objectivity is ranked significantly lower. This finding aligns with the work of Liao et al. (2022), which poses a similar question but focuses on a pool of experts.
Table 12 summarizes the results for Q8 and Q9. According to user feedback, Q1 is deemed the most important question. Notably, there is a strong correlation between the responses to Q8 and Q9: the questions perceived as easiest to answer are also regarded as the most important. Interestingly, there is a low correlation between the importance assigned to each question and the axioms they represent, highlighting a distinction between the perception of an axiom and the execution of the associated task. Furthermore, both Q8 and Q9 reveal a separation in ranking among Q1 to Q4 and Q5 and Q6, which resonates with the dataset analysis discussed in Section B.2.
@@
Appendix C PASTA-metric
C.1 Implementation details
C.1.1 Aggregation of the votes.
[…]
C.1.2 Evaluation metrics
@@
C.2 Ablation studies
C.2.1 Aggregation of the votes
Given the subjective nature of the annotations and the presence of multiple responses to the same question (five answers per question), we explored different methods for determining the ground truth. In Table 13, we tested how PASTA-metric training is affected when using the mean, mode, or median as the ground truth. Since this parameter significantly impacts the dispersion of the samples, it is not surprising that the results vary, particularly when using the mean. However, in all our experiments, we opted to use the mode, as phenomena of high non-consensus were observed (see Appendix B.2).
Label Type MSE QWK SCC
Mean 0.51 $\pm$ 0.03 0.53 $\pm$ 0.04 0.27 $\pm$ 0.27
Mode 0.84 $\pm$ 0.06 0.56 $\pm$ 0.05 0.28 $\pm$ 0.29
Median 0.73 $\pm$ 0.06 0.55 $\pm$ 0.06 0.28 $\pm$ 0.28
C.2.2 Losses
C.3 Loss Functions
Here we define the tree losses used to train our PASTA-model.
The Cosine Similarity Loss measures the cosine similarity between the predicted explanations $\hat{m}_{k}$ and the ground truth explanations $m_{k}$, ensuring alignment in their direction:
@@ C.3 Loss Functions
C.3.1 Explanation embeddings

### v2

@@ B.3 Dataset analysis
To further explore annotator preferences, we identified the top-12 XAI techniques selected by each annotator and visualized the results in the histogram shown in Figure 16. Unlike Table 11, this analysis focuses on the top-12 techniques per annotator, removing the influence of votes among the top-12 techniques to reduce noise and better capture annotator preferences. From this figure, we observe that classical methods such as LIME and SHAP stand out as one of the most frequently preferred. This suggests a strong preference for well-established saliency-based methods. Additionally, there is a notable inclination towards methods that probe the model’s reaction to input perturbations. A distinct aspect is the inclusion of B-cos, which generates explanations through the incorporation of a dedicated layer, offering a unique mechanism compared to other perturbation-based techniques. Of the 11 most popular techniques among annotators, only two are based on CBMs, indicating a general preference for saliency maps over concept-based explanations.
@@
Next, we aggregated the scores using majority voting and calculated QWC scores to measure the agreement between individual annotators and the aggregated score. We further analyzed the QWC scores by gender and age groups to assess any systematic differences in interpretation. As shown in Figure 17, the kappa scores indicate that there is generally consistent agreement across different age and sex groups, although older annotators show slightly less consistency. This highlights that while demographic factors may introduce some variation, they do not substantially impact the overall interpretability evaluation. Notable exceptions are observed for Q3 and Q4, which address the notions of objectivity and clarity. The decrease in agreement can be attributed to the increased subjectivity required by these questions compared to others.
Refer to caption Refer to caption
In Table 11, we present the average mode of votes for each XAI technique. The primary observation is the challenge in discerning clear differences among XAI methods, largely attributable to their high sensitivity to the specific backbones to which they are applied. Regarding the differences among questions, we note significant similarities between Q1 and Q2, as well as between Q3 and Q4. It is particularly noteworthy that raters globally assessed explanations as clearer and more objective, with mean scores of 3.46 for Q3 and 3.44 for Q4, compared to their faithfulness to the model, which had mean scores of 2.83 for Q1 and 2.82 for Q2. Beyond confirming the multifaceted nature of human assessment, this underscores the difficulties encountered by current XAI methods in conveying the behavior of the model, rather than merely creating clear methods for displaying explanations.
XAI Technique Q1 Q2 Q3 Q4 Q5 Q6
GradCAM (ResNet50) 3.22 3.22 3.67 3.66 2.22 4.65
GradCAM (ViT-B) 2.63 2.64 3.41 3.41 2.53 4.62
GradCAM (CLIP-zero-shot) 2.40 2.40 3.30 3.29 2.63 4.79
LIME (ResNet50) 2.51 2.49 3.21 3.20 3.78 4.51
LIME (ViT-B) 3.00 2.98 3.55 3.53 3.23 4.53
LIME (CLIP-zero-shot) 2.75 2.74 3.35 3.33 3.61 4.53
SHAP (ResNet50) 3.00 2.99 3.50 3.49 2.64 4.70
SHAP (ViT-B) 3.10 3.09 3.58 3.55 2.57 4.70
SHAP (CLIP-zero-shot) 3.14 3.13 3.55 3.52 2.64 4.68
AblationCAM (ResNet50) 2.71 2.70 3.46 3.46 3.07 4.17
AblationCAM (ViT-B) 2.29 2.28 3.21 3.21 2.48 4.66
AblationCAM (CLIP-zero-shot) 2.03 2.03 2.99 3.00 2.34 4.64
EigenCAM (ResNet50) 3.12 3.11 3.65 3.63 2.26 4.25
EigenCAM (ViT-B) 3.13 3.13 3.65 3.63 2.27 4.29
EigenCAM (CLIP-zero-shot) 3.27 3.28 3.81 3.76 1.98 4.11
EigenGradCAM (ResNet50) 3.09 3.09 3.61 3.60 2.13 4.32
EigenGradCAM (ViT-B) 2.00 2.00 2.90 2.90 2.98 4.69
EigenGradCAM (CLIP-zero-shot) 2.73 2.73 3.45 3.43 2.62 4.58
FullGrad (ResNet50) 3.50 3.49 3.86 3.80 1.98 4.03
FullGrad (ViT-B) 2.28 2.27 3.16 3.16 3.13 4.64
GradCAMPlusPlus (ResNet50) 3.26 3.26 3.73 3.69 2.20 4.38
GradCAMPlusPlus (ViT-B) 2.58 2.58 3.33 3.31 3.10 4.69
GradCAMPlusPlus (CLIP-zero-shot) 2.31 2.31 3.27 3.26 2.78 4.79
GradCAMElementWise (ResNet50) 3.23 3.22 3.69 3.67 2.19 4.32
GradCAMElementWise (ViT-B) 2.12 2.12 2.95 2.97 2.69 4.65
GradCAMElementWise (CLIP-zero-shot) 3.10 3.10 3.64 3.61 2.25 4.37
HiResCAM (ResNet50) 3.28 3.27 3.72 3.69 2.21 4.60
HiResCAM (ViT-B) 2.03 2.03 2.98 2.98 2.93 4.80
HiResCAM (CLIP-zero-shot) 2.38 2.38 3.21 3.21 2.73 4.80
ScoreCAM (ResNet50) 3.32 3.32 3.73 3.70 2.32 4.50
ScoreCAM (ViT-B) 3.19 3.18 3.73 3.68 2.44 4.66
XGradCAM (ResNet50) 3.26 3.27 3.71 3.69 2.10 4.63
XGradCAM (ViT-B) 2.37 2.37 3.29 3.28 3.78 4.65
XGradCAM (CLIP-zero-shot) 2.59 2.59 3.46 3.45 3.88 4.70
DeepFeatureFactorization (ResNet50) 3.41 3.41 3.86 3.82 2.22 4.18
DeepFeatureFactorization (ViT-B) 3.03 3.02 3.72 3.68 2.18 4.26
DeepFeatureFactorization (CLIP-zero-shot) 3.22 3.22 3.78 3.72 2.02 4.07
CLIP-QDA-sample 2.20 2.19 3.11 3.09 2.72 4.70
CLIP-Linear-sample 3.10 3.08 3.48 3.45 1.68 4.46
LIME_CBM (CLIP-QDA) 2.69 2.69 3.38 3.36 3.64 4.59
SHAP_CBM (CLIP-QDA) 3.09 3.08 3.53 3.52 2.82 4.54
LIME_CBM (CBM-classifier-logistic) 2.90 2.88 3.38 3.35 2.77 4.49
SHAP_CBM (CBM-classifier-logistic) 3.08 3.06 3.46 3.43 2.43 4.30
Xnesyl-Linear 2.36 2.34 3.22 3.19 2.45 3.83
BCos (ResNet50-BCos) 2.86 2.85 3.49 3.46 2.30 3.84
RISE-CBM (CBM-classifier-logistic) 3.28 3.27 3.66 3.62 2.22 4.35
Appendix C C. PASTA-dataset: Extra questions
To enhance the study, we solicited responses from a subset of annotators (15 annotators) to address additional questions. These questions did not directly involve generated explanations. Instead, the questions focused on their overall perceptions of XAI and how they personally approach explanations. Consequently, this section aims to take a broader perspective and reflect more comprehensively from a user study viewpoint.
C.1 Image labeling
For each image, participants had to explain what led them to classify the displayed image as the model’s prediction. Participants responded openly using a text form.
A first set of questions aims at having annotators establish a baseline, i.e., by interpreting and explaining what makes an image recognizable as a specific object or class. The process is constituted of theses two questions:
Similarly, they were asked to describe the elements of the image that helped them make the decision to classify the image as the model did. These two questions (Q0.1 and Q0.2) are used to establish a baseline for interpreting what makes an image recognizable as a specific object or class, and what are the salient features of the images that would explain this choice.
@@ C.1 Image labeling
We investigated potential biases in the annotations themselves by examining the differences in how annotators approached CBM-based and saliency-based explanations. For CBM explanations, we focused on the text written by annotators in response to question Q0.1, assessing whether annotators preferred explanations that closely resembled their own textual responses. To quantify this, we transformed CBM explanations into text by concatenating the top concepts used in the explanation and calculated the BLEU (Papineni et al. 2002) and ROUGE scores (Lin 2004) between these explanations and the annotators’ text responses. As shown in Figure 18, the ROUGE score reveals a slight correlation between the explanations and the annotators’ expectations for questions Q1, Q2, Q3, and Q4. This suggests that annotators are inclined to favor explanations that align with their preconceived notions, potentially introducing a bias toward consistency with their initial answers.
@@
Figure 18: Correlation between the various questions and key metrics for saliency and CBM explanations. For CBM, the criteria used are the BLEU (Papineni et al. 2002) and ROUGE scores (Lin 2004) scores between the explanation and the text from question Q0.1. For saliency maps, the metrics include the total pixel sum (SUM_all), the sum of pixels within the annotator-provided bounding box (SUM_pos) from question Q0.2, the sum of pixels outside the bounding box (SUM_neg), and the entropy of the saliency map (Entropy).
C.2 Desiratas ranking
After the annotation process is completed, evaluators are presented with a set of questions assessing their stance on the questions asked.
@@ C.2 Desiratas ranking
  • Q7: Could you rank the different qualities of the explanation in order of importance? Expectedness, Trustworthiness, Complexity, Robustness, Objectivity.
@@
The results for Q7 are presented in Table 12. The data indicate that evaluators place a higher value on Trustworthiness and Complexity, while Objectivity is ranked significantly lower. This finding aligns with the work of Liao et al. (2022), who asked a similar question but focused on a pool of experts.
Table 13 summarizes the results for Q8 and Q9. According to user feedback, Q1 is deemed the most important question. Notably, there is a strong correlation between the responses to Q8 and Q9: the questions perceived as easiest to answer are also regarded as the most important. Interestingly, there is a low correlation between the importance assigned to each question and the axioms they represent, highlighting a distinction between the perception of an axiom and the execution of the associated task. Furthermore, both Q8 and Q9 reveal a separation in ranking among Q1 to Q4 and Q5 and Q6, which resonates with the dataset analysis.
@@
Appendix D D. PASTA-score: Implementation details
D.1 Technical information
Hyperparameters
In all experiments, we employed the Adam optimizer (Kingma and Ba 2017) with a batch size of 256, training for 600 epochs at a learning rate of $2\times 10^{-6}$. Additionally, we configured the parameters as follows: $\alpha$ was set to 1, $\beta$ to $0.001$, and $\gamma$ to $0.01$. We also applied a weight decay of $1\times 10^{-6}$. The model utilized a hidden layer size of [512, 64]. For concept-based explanations, we used the top 15 concepts, with no template. For Q3 to Q6, we excluded a part of the annotations related to COCO, due to a presence of outliers.
Regarding the division between the training, validation, and test sets, we imposed restrictions on the selected samples to ensure no overlap in XAI methods or images on which the explanations are based, thereby preventing overfitting. To achieve this, we proceeded as follows. First, we assigned an ID to each of the images (IMG_ID) and each of the XAI methods (XAI_ID). As mentioned in the dataset specifications, $\text{IMG\_ID}\in\llbracket 1~;~1000\rrbracket$ and $\text{XAI\_ID}\in\llbracket 1~;~46\rrbracket$. Secondly, given these sets and a seed, we randomly selected 70% of the XAI_ID and 70% of the IMG_ID. Thirdly, we included in the training set only those samples that presented both a valid XAI_ID and IMG_ID ($\neg\left((\text{XAI\_ID\_train}=\text{XAI\_ID\_test})\lor(\text{IMG\_ID\_train}=\text{IMG\_ID\_test})\right)$). We followed the same procedure with the remaining set to construct the validation and test sets. All remaining samples were discarded for the run.
Evaluation metrics
@@ D.1 Technical information
Loss Functions
Here, we define the three losses used to train our PASTA-model.
The Cosine Similarity Loss measures the cosine similarity between the predicted explanations scores $\hat{m}_{k}$ and the ground truth explanations scores $m_{k}$, ensuring alignment in their direction:
@@
Computational resources
The experiments were conducted using a V100-32GB GPU. The training and inference times are summarized in Table 14. It is important to note that, for both inference and testing, the majority of the computational time is dedicated to precomputing VLM embeddings, as the scoring network itself is relatively lightweight and requires minimal computational resources.
Table 14: Training and Inference Times on GPU and CPU. Results for the $\text{PASTA-score}^{\text{SIGLIP}}$ variant.
Device Training Time (s) Inference Time (s)
GPU 1048.74 36.12
CPU 10035.98 796.41
D.2 Ablation studies
Aggregation of the votes
Formulas
$\text{Mode}(m_{i}^{j})=\underset{a}{\arg\max}\;\text{Count}(m_{i}^{j}[a])\,$, (13)
Given the subjective nature of the annotations and the presence of multiple responses to the same question (five answers per question), we explored different methods for determining the ground truth. In Table 15, we tested how PASTA-score training is affected when using the mean, mode, or median as the ground truth. Since this parameter significantly impacts the dispersion of the samples, it is not surprising that the results vary, particularly when using the mean. However, in all our experiments, we opted to use the mode, as phenomena of high non-consensus were observed.
Label Type MSE QWK SCC
Mode 0.989 $\pm$ 0.113 0.471 $\pm$ 0.055 0.501 $\pm$ 0.052
Mean 0.774 $\pm$ 0.067 0.506 $\pm$ 0.049 0.532 $\pm$ 0.043
Median 0.924 $\pm$ 0.093 0.479 $\pm$ 0.052 0.509 $\pm$ 0.046
Add of label information in the PASTA-score embedding.
In the current iteration of the PASTA-score, information regarding the predicted class is integrated into the embeddings supplied to the scoring network. Specifically, each predicted label is encoded as a one-hot vector and concatenated with the embedding. In this study, we aim to measure the impact of this design choice by comparing it with an alternative approach that utilizes only embeddings as input to the scoring network.
As illustrated in Table 16, the integration of label information exhibits a positive influence, aligning with the hypothesis that such contextual data enhances the comprehension of human behavior patterns affecting ratings. However, the observed improvement is modest. This phenomenon can be attributed to several factors, including the extensive number of labels employed across the datasets (26), which is commensurate with the quantity of unique images. As a result, the inclusion of label information may introduce redundancy or lead to overfitting, thereby influencing the efficiency of the scoring mechanism.
Table 16: Impact of Adding Labels to the Encodings. Each value is the average result on 5 runs with the standard deviation.
Computation MSE QWK SCC
Embeddings 0.987 $\pm$ 0.098 0.469 $\pm$ 0.047 0.497 $\pm$ 0.045
Embeddings + Labels 0.989 $\pm$ 0.113 0.471 $\pm$ 0.055 0.501 $\pm$ 0.052
Scoring functions
In this section, we examine the influence of various scoring network architectures on performance. Specifically, we tested alternatives such as Ridge Regression, Lasso Regression, Support Vector Machines. The results of these experiments are presented in Table 17.
By analyzing the performance of the different scoring functions, we observe that PASTA, implemented with a multi linear perceptron and leveraging the loss functions described in Section Scoring network, achieves superior results in terms of the Quadratic Weighted Kappa score and Spearman Correlation Coefficient. These outcomes highlight its effectiveness in accurately ranking labels. However, SVM exhibit lower Mean Square Error, suggesting better numerical precision in predicting label values. Our primary objective is to develop a robust metric for ranking XAI methods. As such, we place greater emphasis on metrics that assess ranking accuracy. Based on this criterion, the PASTA framework is favored over alternative scoring networks due to its superior performance in rank-oriented evaluations.
Table 17: Impact of the Scoring Function. Each value is the average result on 5 runs with the standard deviation.
Scoring Function MSE QWK SCC
PASTA 0.989 $\pm$ 0.113 0.471 $\pm$ 0.055 0.501 $\pm$ 0.052
SVM 0.965 $\pm$ 0.077 0.463 $\pm$ 0.056 0.501 $\pm$ 0.051
Ridge 1.169 $\pm$ 0.143 0.423 $\pm$ 0.046 0.444 $\pm$ 0.043
Lasso 0.998 $\pm$ 0.075 0.395 $\pm$ 0.048 0.454 $\pm$ 0.052
Explanation embeddings
This section presents several ablation studies that informed the development of our current methodology for embedding explanations. Throughout these experiments, the Siglip model was utilized as the VLM encoder.